OpenStudy (anonymous):

A 2.0m -long, 750g rope pulls a 6.0kg block of ice across a horizontal, frictionless surface. The block accelerates at 1.0m/s2 . A)How much force pulls forward on the block of ice? Express your answer using two significant figures. B)How much force pulls forward on the rope? Express your answer using two significant figures.

OpenStudy (anonymous):

A/ force = mass of block of ice * acceleration = 6*1 B/ total force = total masses of block of ice + rope * acceleration = (0.75+6)*1